# Improving IR-based Bug Localization with Semantics-Driven Query Reduction

### v1

@@ 5.3 Evaluting IQLoc
To determine an optimal length for search queries (N), we use a fixed pre-trained model, the baseline CodeT5 [64] model, during document reranking step while varying the length of queries. From Fig. 11 we see that the HIT@1 increases from 0.32 with 5 keywords to a maximum of 0.455 with a keyword length of 43, representing a $\approx$42% increase. Such a trend is consistent across other metrics, with improvements of 27.64%, 37.61%, 34.61%, and 31.52% for MAP, MRR, HIT@5, and HIT@10, respectively. However, we observe diminishing returns in performance improvement for HIT@K as query length increases, reaching a plateau around the length of 15, where performance improvement slows down. To balance between performance and query specificity, we thus chose a maximum query length of 15 for our experiments.
@@
We also investigate the role of embedding models in our query reformulation step during bug localization, as the quality of reformulated queries directly impacts localization effectiveness. Fig. 12 compares the performance of CodeT5 and CodeBERT in terms of MAP and MRR to determine which model is more effective for generating reformulated queries in our proposed technique. Reformulated queries using CodeBERT achieve a MAP score of 0.306 for a keyword length of 5, which is 27.90% lower than CodeT5’s score of 0.392. Although increasing the query length reduces the performance gap, CodeBERT does not surpass CodeT5. For instance, at a keyword length of 43, CodeBERT still performs 3.71% lower than CodeT5. A similar trend is observed for MRR (Fig. 12b), where CodeBERT lags by 27.35% at a keyword length of 5. Based on these performance differences, we selected CodeT5 for embedding generation in our keyword extraction module, as it leads to more effective query reformulations and thus enhances the overall bug localization performance.
We also wanted to investigate if domain-specific embedding helps in constructing queries. Recognizing the importance of domain-specific pre-training to capture nuanced language features [44], we pre-trained the CodeT5-small [64] model on a comprehensive dataset containing bug reports (see Section 4.5.1 for details). This pre-training aimed to equip the model with a deeper understanding of software bugs from Java-based systems. Our findings indicate that the pre-trained model consistently outperforms the baseline model across all metrics, achieving superior performance even with fewer keywords. From Fig. 13, we see that the MAP improves from 0.39 to 0.42, representing a 9.68% increase over the baseline model using 5 keywords. Similarly, it increases from 0.45 to 0.48 with 15 keywords, indicating a 6.98% improvement. This trend holds true across various query lengths experimented with, ranging from 5 to 43 for all metrics. Therefore, we choose the CodeT5 model pre-trained with bug reports for our technique - IQLoc.
@@
Answering RQ3: Performance of the Cross-Encoder Model in Determining Buggy Source Code Based on Program Semantics
In this section, we investigate the effectiveness of the cross-encoder model (adopted by IQLoc) in identifying buggy code segments based on their program semantics. Transformer-based cross-encoder models are designed to output probabilistic confidence scores ranging from 0 to 1. We fine-tuned our model to classify code segments as either buggy (i.e., 1) or non-buggy (i.e., 0) against on a given bug report. Fig. 14 illustrates the performance of our adopted cross-encoder model with various configurations.
For the random split evaluation set (Fig. 14a), at threshold 0, the model achieves the lowest accuracy of 19.5%, with the lowest precision and F1 scores and the highest recall. This occurs because, at this threshold, the model classifies all instances as positive (a.k.a., buggy), capturing all true positives (TP) but also misclassifying all negative cases as false positives (FP). Since recall is calculated as $TP/(TP+FN)$ and $FN=0$ in this scenario, recall remains at 1 despite poor precision or accuracy. Once the threshold increases above 0 (e.g., 0.1), performance stabilizes, reaching an accuracy of 91.4%, with precision, recall, and F1 scores of 81.3%, 73.8%, and 75.82%, respectively. However, at threshold 1, the opposite effect occurs— the model classifies all instances as negative, correctly identifying all negative cases (TN) but misclassifying all positive cases (TP), leading to a recall of 0. Since our evaluation dataset consists of four times more negative cases than positive ones (discussed in Section 5.1), the accuracy at threshold 1 is approximately 80%. It is noteworthy that these cross-encoder results for the random split evaluation set are averaged over five independent runs.
In the time-wise split scenario (Fig. 14b), our cross-encoder model exhibits similar trends for thresholds 0 and 1. However, once the threshold surpasses 0.1, the model stabilizes and demonstrates slightly improved performance across all metrics. Starting with accuracy, the model shows a consistent upward trend, reaching $\approx$93%, which is a 1.9% improvement over the random split dataset scenario. Similarly, precision, recall, and F1 scores remain relatively stable, hovering around 88.7%, 74.9%, and 81.2%, respectively.
@@
Answering RQ4: Comparison with the Baseline Techniques
@@
In this section, we compare our proposed technique IQLoc against existing bug localization techniques in terms of various evaluation metrics. In particular, we compare IQLoc with five baseline techniques – Baseline Elasticsearch, BLUiR [19], Blizzard [11], DNNLoc [17], and RLocator [31].
@@
To replicate the baseline Elasticsearch technique, we indexed all source documents of a subject system’s repository and capture the pre-processed bug reports as queries. Then, we execute the queries with the Elasticsearch engine [42], which retrieves the relevant source documents based on the BM25 algorithm [65] and Boolean query [66]. We use the default values for the parameters - k and b - in the BM25 algorithm.
@@
BLUiR [19] employs a structured Information Retrieval approach where it leverages the structural components of both source code documents and bug reports for effective localization, which helps avoid spurious matching. To achieve this, BLUiR collects four code elements (i.e., class names, method names, variable names, and comments) from each source code document and two textual elements from each bug report (i.e., bug title and description). Then, it conducts eight searches to compute suspiciousness score for each code-text pair and then combines scores to calculate the overall suspiciousness score for a source document. For the experiments, we collected BLUiR’s replication package from the Bench4BL repository [23] and adapted it for version-based replication. It uses Indri [67], a TF-IDF [68]-based search engine, as its backend for experiments, which has become obsolete recently. Thus, we chose to replicate the technique using Apache Lucene [63] with a TF-IDF-based scoring algorithm. It should be noted that we adopted the formula suggested by Saha et al.[19] for calculating the TF and IDF metrics, which were then fed to the TF-IDF-based retrieval algorithm of Apache Lucene.
@@
Blizzard [11] is an IR-based bug localization technique that leverages contextual information from bug reports for query construction. It categorizes bug reports into three types (i.e., NL-Natural Language, PE-Program Element, and ST-Stack Trace), employs three separate graph-based techniques to construct queries from them, and then retrieves buggy source documents from the corpus by executing the queries with the Apache Lucene engine [63]. For the replication, we collected and adapted the replication package hosted at GitHub [69] by the original authors.
@@
DNNLoc [17] is the first technique for bug localization that combines Deep Learning and Information Retrieval. It uses several features – bug report-source code similarity (rVSM score [18]), class name similarity, collaborative filtering, bug report recency, and bug report frequency to train a deep learning model and then uses the model to localize the bugs. To replicate DNNLoc, we trained appropriate models by extracting features from our training sets, following the authors’ suggestions. During bug localization, we used these trained models to predict suspiciousness scores of source code documents and to rank them.
@@
RLocator [31] is a recent IR-based technique for bug localization that incorporates reinforcement learning [70], modeling the localization task as a Markov Decision Process (MDP) [71]. Before applying reinforcement learning (RL), it retrieves source documents from Elasticsearch based on the bug report and filters suspicious documents using an XGBoost [72] model. It then employs an RL agent based on an actor-critic framework [73] that attempts to rank relevant source documents at the top positions, with MAP or MRR as reward signals. We collected the replication package of RLocator from Zenodo [74] and adapted it for our study on version-based bug localization while maintaining the original specifications of the authors.
@@
From Table 10a (random split dataset), we see that existing techniques such as baseline Elasticsearch achieve a MAP score of 0.457, while BLUiR, Blizzard, and RLocator achieve a comparative MAP score from 0.470 to 0.480, indicating $\approx$2.71%-4.89% improvement. On the other hand, DNNLoc achieves a much lower MAP score of 0.311. IQLoc, on the other hand, has a MAP score of 0.493, which is 2.71% to 58.52% better than the baseline scores. This shows that relevant documents rank higher than irrelevant ones compared to other baseline techniques. Similar improvements are observed for MRR and HIT@1, with increases of up to 61.49% and 69.88%, respectively, indicating promising performance. Although IQLoc exhibits a marginal decrease in HIT@5 and HIT@10 compared to BLUiR, Blizzard, and Rlocator, it demonstrates an improvement of up to 55.53% and 41.9% over the baseline Elasticsearch and DNNLoc measures.
@@
In Table 10b (time-wise split dataset), IQLoc’s performance also surpasses those of baseline techniques. While Elasticsearch achieves a MAP score of 0.471, BLUiR, Blizzard, and RLocator achieve comparative MAP scores between 0.494 and 0.508; DNNLoc’s score is much lower at 0.324. In contrast, IQLoc achieves a MAP score of 0.520, representing improvements ranging from 2.36% to 60.49% over the baseline measures. Similar improvements are observed for MRR and HIT@1, ranging from 2.59%-64.58%, and 6.15%-100.9%, respectively. However, like the randomly split data (Table 10a), IQLoc experiences a marginal drop in HIT@5 and HIT@10 compared to BLUiR and Blizzard. Nonetheless, it demonstrates improvements of up to 40.25% and 24.37% over the baseline Elasticsearch and DNNLoc measures in these metrics.
Refer to caption Refer to caption
(a) Random Split
Refer to caption Refer to caption
(b) Time-wise Split
@@
Figure 15: Comparison of Baseline Techniques in Localizing Different Types of Bugs
Refer to caption Refer to caption
(a) Random Split
Refer to caption Refer to caption
(b) Time-wise Split
@@
Figure 16: Performance of Different Techniques on Different Subject Systems
@@
We also demonstrate the effectiveness of IQLoc in localizing bugs from different types of bug reports compared to baseline techniques. Fig. 15 presents the performance of various methods in localizing bugs from the reports with Stack Trace (ST), Program Elements (PE), and Natural Language (NL) [11]. For the random split test set (Fig. 15a), Elasticsearch, BLUiR, and Blizzard show similar MAP scores (0.54–0.58) for ST bug reports, while RLocator performs slightly better at 0.62. IQLoc outperforms all, achieving 0.66 MAP, a 6.42% improvement over RLocator. In the case of bug reports with program elements (PE), most techniques score between 0.53 and 0.56, except Elasticsearch, which lags at 0.49, while IQLoc leads with 0.56 MAP. DNNLoc consistently underperforms, scoring 0.36 for ST and 0.33 for PE. For natural language-only (NL) bug reports, all techniques struggle, with IQLoc performing comparably. A similar trend is observed in MRR, where IQLoc improves by 6.15%–86.49% for ST and 1.72%–126.92% for PE. In the case of bug reports containing only natural language texts, MRR shows an improvement reaching $\approx$ 64%.
@@
A similar pattern was observed in the time-wise split test set (Fig. 15b), where IQLoc achieves 7.81%–91.67% improvement in ST and 3.63%–72.73% in PE in terms of MAP. For MRR, IQLoc shows up to 80.48% improvement in bug reports with stack traces and 66.67% improvement in those with program elements. However, for bug reports containing only natural language, all techniques perform similarly, except for DNNLoc, which remains the lowest-performing model.
@@
We further analyze each techniques performance across 42 subject systems and compare their MAP and MRR using box plots (Fig. 16). For the random split test set (Fig. 16a), IQLoc achieves a higher median MAP than that of all competitors except BLUiR. However, our technique outperforms others in terms of median MRR.
@@
For the time-wise split test set (Fig. 16b), IQLoc outperforms all techniques with a higher median value and a more compact interquartile range (IQR) for both MAP and MRR, suggesting lower variability in performance. While BLUiR also shows a compact IQR, its median value is lower. DNNLoc, despite being less variable, consistently performs poorly across all subject systems, reinforcing its less effectiveness. Overall, IQLoc demonstrates superior performance in localizing bugs across different subject systems with a higher median measure and a reduced variability in performance.
@@
IQLoc vs p-value Effect-Size (Cliff’s $\delta$)
Elasticsearch 0.0098 Large (0.71)
BLUiR 0.0176 Large (0.58)
Blizzard 0.0209 Medium (0.49)
DNNLoc 0.0063 Very Large (0.78)
RLocator 0.0228 Medium (0.47)
IQLoc’s performance improvements over the baseline measures demonstrate its effectiveness across both dataset splits. To further validate this efficacy, we conducted Wilcoxon signed-rank test [75], a non-parametric statistical test. We use the SciPy library [76] to compute statistical significance of Precision@10 where we capture all buggy documents matching the ground truths within top-10 positions. For our significance analysis, we computed Precision@10 for each of the 1,501 samples across 42 subject systems in our test set, comparing IQLoc against multiple baseline approaches. During the test, we used the alternative=‘less’ parameter (i.e., baseline values are less than IQLoc values) to determine if IQLoc consistently achieved higher precision scores within the top-10 results compared to the baseline. In the case of the time-wise split dataset, IQLoc demonstrated p-values ranging from 0.0063 to 0.0228 when compared against baseline techniques (Table 11) for Precision@10. These p-values are below the significance threshold of 0.05, indicating statistical significance. Moreover, IQLoc exhibited effect sizes ranging from medium to very large, as measured by Cliff’s $\delta$ [77], further explaining the extent of differences. Note that we avoid measuring statistical significance for the randomly split dataset, as we perform five independent runs and average the results for our experiment. Given the evidence above, the null hypothesis is rejected, and IQLoc’s performance is found to be significantly higher than the baseline measures.
RQ4 Summary: IQLoc outperforms baseline techniques significantly, with MAP performance improvements of 58.52% for the random split dataset and 60.49% for the time-wise split dataset. This improvement stems from its ability to localize various types of bugs across different subject systems. Additionally, statistical significance tests confirm the superiority of our technique over baseline techniques with medium to large effect sizes.

### v2

@@ 5.3.2 Answering RQ2: Impact of Query Length and Embedding Models on IQLoc’s Performance
To determine an optimal length for search queries (N), we use a fixed pre-trained model, the baseline CodeT5 [wang2021codet5] model, during document reranking step while varying the length of queries. From Fig. 11 we see that the HIT@1 increases from 0.32 with 5 keywords to a maximum of 0.455 with a keyword length of 43, representing a $\approx$42% increase. Such a trend is consistent across other metrics, with improvements of 27.64%, 37.61%, 34.61%, and 31.52% for MAP, MRR, HIT@5, and HIT@10, respectively. However, we observe diminishing returns in performance improvement for HIT@K as query length increases, reaching a plateau around the length of 15, where performance improvement slows down. To balance between performance and query specificity, we thus chose a maximum query length of 15 for our experiments.
@@
We also investigate the role of embedding models in our query reformulation step during bug localization, as the quality of reformulated queries directly impacts localization effectiveness. Fig. 12 compares the performance of CodeT5 and CodeBERT in terms of MAP and MRR to determine which model is more effective for generating reformulated queries in our proposed technique. Reformulated queries using CodeBERT achieve a MAP score of 0.306 for a keyword length of 5, which is 27.90% lower than CodeT5’s score of 0.392. Although increasing the query length reduces the performance gap, CodeBERT does not surpass CodeT5. For instance, at a keyword length of 43, CodeBERT still performs 3.71% lower than CodeT5. A similar trend is observed for MRR (Fig. 12(b)), where CodeBERT lags by 27.35% at a keyword length of 5. Based on these performance differences, we selected CodeT5 for embedding generation in our keyword extraction module, as it leads to more effective query reformulations and thus enhances the overall bug localization performance.
We also wanted to investigate if domain-specific embedding helps in constructing queries. Recognizing the importance of domain-specific pre-training to capture nuanced language features [sun2020sifrank], we pre-trained the CodeT5-small [wang2021codet5] model on a comprehensive dataset containing bug reports (see Section 4.5.1 for details). This pre-training aimed to equip the model with a deeper understanding of software bugs from Java-based systems. Our findings indicate that the pre-trained model consistently outperforms the baseline model across all metrics, achieving superior performance even with fewer keywords. From Fig. 13, we see that the MAP improves from 0.39 to 0.42, representing a 9.68% increase over the baseline model using 5 keywords. Similarly, it increases from 0.45 to 0.48 with 15 keywords, indicating a 6.98% improvement. This trend holds true across various query lengths experimented with, ranging from 5 to 43 for all metrics. Therefore, we choose the CodeT5 model pre-trained with bug reports for our technique - IQLoc.
@@
5.3.3 Answering RQ3: Performance of the Cross-Encoder Model in Determining Buggy Source Code Based on Program Semantics
In this section, we investigate the effectiveness of the cross-encoder model (adopted by IQLoc) in identifying buggy code segments based on their program semantics. Transformer-based cross-encoder models are designed to output probabilistic confidence scores ranging from 0 to 1. We fine-tuned our model to classify code segments as either buggy (i.e., 1) or non-buggy (i.e., 0) against on a given bug report. The model jointly encodes both inputs (i.e., code and bug reports) and learns their contextual relationships during fine-tuning. Fig. 14 demonstrates the performance of our adopted cross-encoder model under various configurations, highlighting how changes in threshold affect its ability to leverage learned lexical and semantic features from the bug reports and code segments.
For the random split evaluation set (Fig. 14(a)), at threshold 0, the model achieves the lowest accuracy of 19.5%, with the lowest precision and F1 scores and the highest recall. This occurs because, at this threshold, the model classifies all instances as positive (a.k.a., buggy), capturing all true positives (TP) but also misclassifying all negative cases as false positives (FP). Since recall is calculated as $TP/(TP+FN)$ and $FN=0$ in this scenario, recall remains at 1 despite poor precision or accuracy. Once the threshold increases above 0 (e.g., 0.1), performance stabilizes, reaching an accuracy of 91.4%, with precision, recall, and F1 scores of 81.3%, 73.8%, and 75.82%, respectively. However, at threshold 1, the opposite effect occurs— the model classifies all instances as negative, correctly identifying all negative cases (TN) but misclassifying all positive cases (TP), leading to a recall of 0. Since our evaluation dataset consists of four times more negative cases than positive ones (discussed in Section 5.1), the accuracy at threshold 1 is approximately 80%. It is noteworthy that these cross-encoder results for the random split evaluation set are averaged over five independent runs.
In the time-wise split scenario (Fig. 14(b)), our cross-encoder model exhibits similar trends for thresholds 0 and 1. However, once the threshold surpasses 0.1, the model stabilizes and demonstrates slightly improved performance across all metrics. Starting with accuracy, the model shows a consistent upward trend, reaching $\approx$93%, which is a 1.9% improvement over the random split dataset scenario. Similarly, precision, recall, and F1 scores remain relatively stable, hovering around 88.7%, 74.9%, and 81.2%, respectively.
@@
5.3.4 Answering RQ4: Comparison with the Baseline Techniques
@@
In this section, we compare our proposed technique IQLoc against existing bug localization techniques in terms of various evaluation metrics. In particular, we compare IQLoc with nine baseline techniques – Baseline Elasticsearch, BLUiR [saha_bleuir], Blizzard [blizzard], BRTracer [BRTracer], BLIA [bl_code_change_BLIA], LRBL [LRBL_learning2rank], LLmiRQ [LLmiRQ], DNNLoc [dnnloc_ir], and RLocator [RLocator].
@@
To replicate the baseline Elasticsearch technique, we indexed all source documents of a subject system’s repository and capture the pre-processed bug reports as queries. Then, we execute the queries with the Elasticsearch engine [elasticsearch], which retrieves the relevant source documents based on the BM25 algorithm [rank-bm25] and Boolean query [boolean_query]. We use the default values for the parameters - k and b - in the BM25 algorithm.
@@
BLUiR [saha_bleuir] employs a structured Information Retrieval approach where it leverages the structural components of both source code documents and bug reports for effective localization, which helps avoid spurious matching. To achieve this, BLUiR collects four code elements (i.e., class names, method names, variable names, and comments) from each source code document and two textual elements from each bug report (i.e., bug title and description). Then, it conducts eight searches to compute suspiciousness score for each code-text pair and then combines scores to calculate the overall suspiciousness score for a source document. For the experiments, we collected BLUiR’s replication package from the Bench4BL repository [bench4bl] and adapted it for version-based replication. It uses Indri [Indri], a TF-IDF [tf-idf]-based search engine, as its backend for experiments, which has become obsolete recently. Thus, we chose to replicate the technique using Apache Lucene [lucene] with a TF-IDF-based scoring algorithm. It should be noted that we adopted the formula suggested by Saha et al.[saha_bleuir] for calculating the TF and IDF metrics, which were then fed to the TF-IDF-based retrieval algorithm of Apache Lucene.
@@
Blizzard [blizzard] is an IR-based bug localization technique that leverages contextual information from bug reports for query construction. It categorizes bug reports into three types (i.e., NL-Natural Language, PE-Program Element, and ST-Stack Trace), employs three separate graph-based techniques to construct queries from them, and then retrieves buggy source documents from the corpus by executing the queries with the Apache Lucene engine [lucene]. For the replication, we collected and adapted the replication package hosted at GitHub [Blizzard_git_repo] by the original authors.
BRTracer [BRTracer] is an IR-based technique that aims to improve bug localization through document segmentation and stack-trace analysis. Rather than considering each source document as a single unit, it divides the document into equal-sized segments and treats each segment as a corpus document, reducing noise from unrelated parts of large documents. Additionally, it extracts document names from stack traces in bug reports, boosts the ranking of documents appearing in the top 10 stack-trace frames, and assigns a smaller boost score to the remaining related documents. The final ranking combines the segment-based similarity (i.e., rVSM [ir_localization_lda_buglocator]) and the boost scores. We collected the replication package of BRTracer from Bench4BL [bench4bl] and adapted it for our version-based experiments.
BLIA [bl_code_change_BLIA] is another IR-based bug localization technique that combines textual similarity between bug reports and source documents with structured information from code, stack-trace clues, and code change histories. It adopts the revised Vector Space Model (rVSM) for similarity scoring and extends with structured information retrieval [saha_bleuir]. The technique also considers similar past bug reports, boosts documents mentioned in stack traces, and leverages recent commit logs filtered by keywords. These analyses are then integrated with tunable parameters to compute a final suspiciousness score for ranking the source documents. We collected and adapted the BLIA replication package from Bench4BL [bench4bl] for our experiments.
LRBL [LRBL_learning2rank] is a Learning-to-Rank (LtR)-based bug localization technique that employs a data-driven ranking model to localize software bugs. It formulates bug localization as a feature-based learning-to-rank problem, where each bug report–source document pair is represented by six features capturing lexical and API-enriched semantic similarity to the bug report, similarity to previous bug reports associated with the same document (collaborative filtering), the presence of the document’s class name in the bug report, and the document’s historical bug-fixing recency and frequency. Then these pairs are labeled based on historical bug fix data and trained with SVMrank [SVMRank] to rank the suspicious source documents higher in the ranked list. For the experiments, we replicated the technique carefully by following the steps specified by the author.
LLmiRQ [LLmiRQ] is a hybrid technique for bug localization that integrates Large Language Model (LLM)–driven query generation with a supervised Learning-to-Rank (LtR) model. It first classifies each bug report into three categories - program elements (PE), stack trace (ST), and natural language (NL), and uses LLM prompting to reduce search queries for PE and ST reports, and expand queries for NL reports. Using the generated query, it retrieves candidate source documents and trains an SVMrank [SVMRank] model on each (bug report, document) pair with features such as textual similarity, class-name overlap, call-graph relationships, and historical relevance. Then, LLmiRQ employs the trained model to rank potentially buggy documents. We replicated the techniques carefully by following the author’s specifications for our experiment.
@@
DNNLoc [dnnloc_ir] is the first technique for bug localization that combines Deep Learning and Information Retrieval. It uses several features – bug report-source code similarity (rVSM score [ir_localization_lda_buglocator]), class name similarity, collaborative filtering, bug report recency, and bug report frequency to train a deep learning model and then uses the model to localize the bugs. To replicate DNNLoc, we trained appropriate models by extracting features from our training sets, following the authors’ suggestions. During bug localization, we used these trained models to predict suspiciousness scores of source code documents and to rank them.
@@
RLocator [RLocator] is a recent IR-based technique for bug localization that incorporates reinforcement learning [reinforcement_learning], modeling the localization task as a Markov Decision Process (MDP) [markov_decision_reinforcement]. Before applying reinforcement learning (RL), it retrieves source documents from Elasticsearch based on the bug report and filters suspicious documents using an XGBoost [chen2016xgboost] model. It then employs an RL agent based on an actor-critic framework [konda2000actorcritic] that attempts to rank relevant source documents at the top positions, with MAP or MRR as reward signals. We collected the replication package of RLocator from Zenodo [rlocator_replication_zenodo] and adapted it for our study on version-based bug localization while maintaining the original specifications of the authors.
@@
From Table 10(a) (random split dataset), we see that existing techniques such as baseline Elasticsearch achieve a MAP score of 0.457, while BLUiR, Blizzard, and RLocator achieve a comparative MAP score from 0.470 to 0.480, indicating $\approx$2.71%-4.89% improvement. Comparatively, BRTracer achieves a much lower MAP score of 0.246. IQLoc, on the other hand, has a MAP score of 0.493, which is 2.71% to 100.4% higher than the baseline scores. In other words, our technique can rank the relevant documents higher than irrelevant ones, where the other baseline techniques might struggle. Similar improvements are observed for MRR and HIT@1, with increases of up to 61.49% and 69.88%, respectively, indicating promising performance. Although IQLoc exhibits a marginal decrease in HIT@5 and HIT@10 compared to BLUiR, Blizzard, and Rlocator, it demonstrates an improvement of up to 55.52% and 57.18% over the baseline Elasticsearch and DNNLoc, BRTracer, BLIA, LRBL, and LLmiRQ measures.
@@
In Table 10(b) (time-wise split dataset), IQLoc’s performance also surpasses that of baseline techniques. While Elasticsearch achieves a MAP score of 0.471, BLUiR, Blizzard, LRBL, LLmiRQ, and RLocator achieve comparative MAP scores between 0.422 and 0.508; Similar to random split performance, BRTracers, BLIA, and DNNLoc’s scores range from 0.292 to 0.324. In contrast, IQLoc achieves a MAP score of 0.520, representing improvements ranging from 2.36% to 78.08% over the baseline measures. Similar improvements are observed for MRR and HIT@1, ranging from 2.59%-64.58%, and 6.15%-100.9%, respectively. However, like the randomly split case (Table 10(a)), IQLoc experiences a marginal drop in HIT@5 and HIT@10 compared to BLUiR and Blizzard. Nonetheless, it demonstrates improvements of up to 40.25% and 24.36% over the baseline measures in these metrics.
@@
Figure 15: Comparison of Baseline Techniques in Localizing Different Types of Bugs (Random Split)
Figure 16: Comparison of Baseline Techniques in Localizing Different Types of Bugs (Time-wise Split)
@@
Figure 17: Performance of Different Techniques on Different Subject Systems (Random Split)
Figure 18: Performance of Different Techniques on Different Subject Systems (Time-wise Split)
@@
We also demonstrate the effectiveness of IQLoc in localizing bugs for different types of bug reports and compare it with the baseline techniques. Fig. 15 and Fig. 16 present the performance of various techniques on bug reports containing Stack Traces (ST), Program Elements (PE), and Natural Language (NL) [blizzard]. For the random-split test set (Fig. 15), LRBL and LLmiRQ report similar MAP scores ($\sim$0.46) for ST bug reports. Elasticsearch, BLUiR, and BLIZZARD obtain MAP scores in the 0.54-0.58 range, with RLocator achieving a modest improvement at 0.62. IQLoc outperforms all techniques, achieving a MAP 0.66, a 6.42% improvement over RLocator. For PE bug reports, BLUiR, Blizzard, and RLocator score between 0.53 and 0.56, whereas Elasticsearch lags at 0.49; IQLoc again leads with a MAP score of 0.56. BRTracer, BLIA, and DNNLoc consistently underperform, scoring 0.30 to 0.36 for ST and 0.27 to 0.33 for PE. For NL-only bug reports, all techniques struggle, with IQLoc performing comparably. A similar pattern is observed in MRR, where IQLoc improves by 5.69%-85.78% for ST and 1.38%-71.09% for PE. For NL-only bug reports, MRR improvements over the baselines reach $\approx$75%.
@@
A similar pattern is observed in the time-wise split test set (Fig. 16), where IQLoc achieves an 8.43%-118.71% improvement over the baselines for ST and a 2.94%-88.48% improvement for PE in terms of MAP. For MRR, IQLoc shows improvements of up to 79.67% for ST bug reports and 67.02% for those containing program elements. However, for bug reports containing only natural language, all techniques perform similarly, except DNNLoc, BRTracer, and BLIA, which remain the lowest-performing baseline techniques.
@@
We further analyze each technique’s performance across 42 subject systems in our dataset and compare their MAP, MRR and HIT@10 performance using box plots (Fig. 17 and Fig. 18). For the random split test set (Fig. 17), IQLoc achieves a higher median MAP than that of all competitors except BLUiR. However, our technique outperforms all baselines in terms of median MRR. For HIT@10, we observe a slight drop in median performance compared to RLocator, but IQLoc still performs better overall than the other techniques.
@@
For the time-wise split test set (Fig. 18), IQLoc outperforms all techniques with a higher median value and a more compact interquartile range (IQR) for MAP, MRR, and HIT@10, indicating lower variability. Although BLUiR exhibits a compact IQR, its median performance remains lower. BRTracer, BLIA, LRBL, LLmiRQ, and DNNLoc, despite showing low variability, consistently perform poorly across all subject systems, reinforcing their lower effectiveness. Overall, IQLoc demonstrates superior performance in localizing bugs across different subject systems with a higher median measure and a reduced variability in performance.
@@
IQLoc vs p-value (P@10) p-value (RR@10) p-value (HIT@10)
Elasticsearch 0.0098 (Large) 0.0139 (Medium) 0.0198 (Medium)
BLUiR 0.0176 (Large) 0.0296 (Medium) 0.0426 (Small)
Blizzard 0.0209 (Medium) 0.0328 (Medium) 0.0409 (Small)
BRTracer 0.0061 (Very Large) 0.0094 (Very Large) 0.0111 (Large)
BLIA 0.0082 (Very Large) 0.0115 (Large) 0.0202 (Large)
LRBL 0.0132 (Large) 0.0377 (Medium) 0.0294 (Medium)
LLmiRQ 0.0112 (Large) 0.0182 (Large) 0.0112 (Medium)
DNNLoc 0.0063 (Very Large) 0.0091 (Very Large) 0.0103 (Very Large)
RLocator 0.0228 (Medium) 0.0391 (Small) 0.0377 (Small)
P@10 = Precision@10, RR@10 = Reciprocal Rank, = Effect size (Cliff’s $\delta$)
IQLoc’s performance improvements over the baseline measures demonstrate its effectiveness across both dataset splits. To further validate this efficacy, we conducted Wilcoxon signed-rank test [wilcoxon_signed_rank], a non-parametric statistical test. We use the SciPy library [scipy] to compute statistical significance of Precision@10 where we capture all buggy documents matching the ground truths within top-10 positions. For our significance analysis, we computed Precision@10 for each of the 1,501 samples across 42 subject systems in our test set (Section 5.1.1, Table 4), comparing IQLoc against multiple baseline approaches. During the test, we used the alternative=‘less’ parameter (i.e., baseline values are less than IQLoc values) to determine if IQLoc consistently achieved higher precision scores within the top-10 results compared to the baseline. In the case of the time-wise split dataset, IQLoc demonstrated p-values ranging from 0.0061 to 0.0228 when compared against baseline techniques (Table 11) for Precision@10. These p-values are below the significance threshold of 0.05, indicating statistical significance. We observe similar trends for Reciprocal Rank (RR@10) and HIT@10, although the degree of significance varies across metrics. Moreover, IQLoc exhibited effect sizes ranging from medium to very large, as measured by Cliff’s $\delta$ [statistical_significance_mann_whitney_cliff], further explaining the extent of differences in Precision@10. While the effect sizes against RLocator, BLUiR, and Blizzard are smaller in terms of HIT@10, IQLoc still maintains superior performance over other baseline techniques. Note that we avoid measuring statistical significance for the randomly split dataset, as we perform five independent runs and average the results for our experiment. Given the evidence above, the null hypothesis is rejected, and IQLoc’s performance is found to be significantly higher than the baseline measures.
RQ4 Summary: IQLoc outperforms baseline techniques significantly, with MAP improvements of up to 100.4% for the random split dataset and up to 78.08% for the time-wise split dataset. Statistical tests confirm the technique’s superiority across different metrics. This improvement stems from its ability to localize various types of bugs across different subject systems.